OpenStudy (anonymous):

WILL GIVE MEDAL. HELP! If number A is a 2-digit number and its digits are transposed to form number B, then the difference between the larger of the two numbers and the smaller of the two numbers must be divisible by _____.

OpenStudy (anonymous):

who knows? lets try some ... \[64-46=18\]

OpenStudy (anonymous):

divisible by 2, 3,6,9 lets try another \[51-14=36\] hmm lets try some more

OpenStudy (anonymous):

\[54-45=9\] ok maybe nine?

OpenStudy (anonymous):

wanna do it for real?

OpenStudy (anonymous):

try \[10a+b-(10b+a)\] get \[10a-a-10b+b=9a-9b\]